OpenStudy (anonymous):

x^2 +4x-4=0 please explain how to solve this quadratic equation.....

OpenStudy (anonymous):

complete the sqaure would be fastest.

OpenStudy (anonymous):

x^2 +4x+4 -4 -4 =0

OpenStudy (anonymous):

(x+2)^2 -8 =0

OpenStudy (anonymous):

\[(x+2)= +- \sqrt{8}\]

OpenStudy (anonymous):

\[x=-2+-2\sqrt{2}\]

OpenStudy (anonymous):

okay thank you !